OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = –5 when x = 15, find y when x = –1

OpenStudy (anonymous):

a. y=-1/5x:1/5 b. y=-1/3x:2/5 c. y=1/3x:-1/3 d. y=-1/3x:1/3

OpenStudy (anonymous):

Y=-1/3x, y= 1/3

OpenStudy (anonymous):

X=....

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thx